Grade 11Organic Chemistry

calcium carbonate reacts with aques solution of hcl to give cacl2 and co2 according tothe reaction, caco3+2hcl gives cacl2+co2+h20.what is the mass of caco3 requiredto react completely with 25 ml of 0.75M HCL.

Aditya majumder
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
calculation of mass of CaCO3 , 
CaCO3 + 2HCl ------> CaCl2 + CO2 + H2O
here we see that, 
2 mole of HCl reacts with 1 mole of CaCO3.
so, 2 × 36.5g of HCl reacts with 100g of CaCO3.
so, 73g of HCl reacts with 100g of CaCO3.
so, 1g of HCl reacts with 100/73 g of CaCO3.
so, 0.6844 g of HCl reacts with 100 × 0.6844/73g of CaCO3 = 68.44/73 g = 0.9375g 
0.9375g of CaCO3 is required to react completely with 25ml of 0.75M HCl